OpenStudy (anonymous):

PLease help, I cannot figure out the formula to use. An investor invested a total of $900 in two mutual funds. One fund earned 9% profit, and the other fund earned %5 profit. If the investor earned a total of $65, how much was in each fund?

OpenStudy (jdoe0001):

so he invested in 2 mutual funds let's call them "a" and "b" how much was the total amount he invested in both? how much money did he make in both total?

OpenStudy (anonymous):

would it be a+b=65 and 9a+5b=900

OpenStudy (jdoe0001):

close, a + b = 900 0.09a + 0.05b = 65

OpenStudy (jdoe0001):

9% = 9/100 n% = n/100

OpenStudy (jdoe0001):

you can multiply the 2nd equation by 100 to make it 9a + 5b = 6500

OpenStudy (anonymous):

I got them backwards and forgot to put the percents in correctly, thank you so very much :)

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

going to solve will you double check it for me when I am done?

OpenStudy (anonymous):

I am doing something wrong, not sure if it's because I have been doing math so long or what. Math makes me so nervous.

OpenStudy (anonymous):

I still need help with this :(

OpenStudy (anonymous):

Would it be the %9 fund had $45, and the %5 fund had 20??????